Betty is 6 years older than Teresa. Three years ago, Betty's age was four times Teresa's age. Find the age of each girl now?
Answer by HyperBrain(694)

let T+6=Betty's present age
T=Theresa's present age.
Three years ago, Betty's age was four times Teresa's age. 3 years ago, betty was (t+6)-3=t+3 years old while Theresa was t-3 years old.
t+3=4(t-3)
t+3=4t-12
3=3t-12
15=3t
t=5>>theresa's age
betty's age=t+6=5+6=11
Therefore, Betty is 11 years old while Betty is 11 years old.
pastime blogtime
Power up,
HyperBrain!